You have recently finished your MBA at the DarnIt Business School. Naturally, you must purchase a new BMW immediately. The car costs about $21,000. The bank quotes an interest rate of 15% APR for a 72 month loan with a 10 percent down payment. You plan on trading the car in for a new one in two years.

1)What will your monthly payment be?

2)What is the effective annual interest rate be on the loan?

3)What will the loan balance be when you trade in the car?
Answer #1

Loan Amount= 0.90(21,000) = $18.900

Calculating Monthly payment,

Using TVM Calculation,

PMT = [FV = 0, PV = 18,900, N = 72, I = 0.15/12]

PMT = $399.64

EAR = (1 + 0.15/12)^12 - 1

EAR = 16.08%

Calculating loan balance in 2 years,

Using TVM Calculation,

FV =[PV= 18,900, PMT = -399.64, N = 24, I = 0.15/12]

FV = $14,359.70